# On the magnetic perturbation theory for Chern insulators

### v1

@@ 1.1. Long-range magnetic perturbations
We begin by describing two simple situations that highlight the singular character of magnetic perturbations. In dimension $d=2$, the free Laplacian $-\Delta$ is a selfadjoint operator on $H^{2}({\mathbb{R}}^{2})$ and its spectrum is purely absolutely continuous and coincides with $[0,\infty)$. When we allow a non-zero constant magnetic field oriented perpendicular to the plane, we have to consider the operator $(-\mathrm{i}\nabla-bA_{L})^{2}$, $b>0$, where $A_{L}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ is the magnetic vector potential in the-called symmetric gauge $A_{L}({\bf x})=\frac{1}{2}(-x_{2},x_{1})$. This is the well-known Landau Hamiltonian, which is selfadjoint on the magnetic Sobolev space $H^{2}_{A_{L}}({\mathbb{R}}^{2})$, given by
$H^{2}_{A_{L}}({\mathbb{R}}^{2})=\overline{\left\{f\in C^{\infty}_{0}({\mathbb{R}}^{2})\,|\,(-\mathrm{i}\nabla-bA_{L})^{2}f\in L^{2}({\mathbb{R}}^{2})\right\}}^{\|\cdot\|_{D_{A_{L}}}}$ (1.1)
@@
whose coefficients grow polynomially at infinity. This strongly suggests that the “magnetic perturbation” $W_{b}$ is neither form, nor operator bounded with respect to the free Laplacian. In general it can be proved that the perturbed magnetic Hamiltonian fails to converge in the norm resolvent sense to the unperturbed non-magnetic Hamiltonian when the constant magnetic field goes to zero, see [1, Theorem 6.3].
It turns out that most of the singular behavior induced by long-range magnetic fields is contained in some highly oscillating unimodular phase factors. In the early 2000s, G. Nenciu and the first Author developed a general method which deals with these singular phases in a systematic way [13, 37]. This framework is nowadays called gauge covariant magnetic perturbation theory. The main purpose of this paper is to outline the main ideas behind this method, together with some physically and mathematically relevant applications.
@@ 1.2. Mathematical framework and results
$\sup_{{\bf x}\in{\mathbb{R}}^{2}}|\partial^{\alpha}b({\bf x})|\leq b\,b_{\alpha},\quad\alpha\in\mathbb{N}^{2},\quad|\alpha|\leq 1,\quad b_{\alpha}<\infty,\quad 0\leq b\leq 1.$ (1.2)
Since $V$ is infinitesimally bounded with respect to the free Laplacian[40, Theorem XIII.96] and the magnetic potentials are smooth, by using the diamagnetic inequality [1, Theorem 2.3] and the standard theory of Schrödinger operators we get that $H_{b}$ is essentially selfadjoint on $C^{\infty}_{0}({\mathbb{R}}^{2})$ and its domain of selfadjointness is the magnetic Sobolev space $H^{2}_{A}({\mathbb{R}}^{2})$, see (1.1) for the definition of the second magnetic Sobolev space.
@@
$C(P)=\lim_{L\to\infty}\frac{2\pi\mathrm{i}}{4L^{2}}\int_{[-L,L]^{2}}\left(P[[P,X_{1}],[P,X_{2}]]\right)({\bf x};{\bf x})\,\mathrm{d}{\bf x}$ (1.3)
exists and it is an integer. The quantity $C(P)$ is called the Chern character of the projection $P$, which can also be interpreted as the trace per unit surface of the double commutator $\mathrm{i}\left(P[[P,X_{1}],[P,X_{2}]]\right)$. The name “Chern” orginates from the fact that (1.3) coincides with the Chern number of the Bloch bundle generated by $P$, whenever $H_{b}$ commutes with a unitary representation of the group ${\mathbb{Z}}^{d}$ provided by magnetic translations [39]. If $P$ is the Fermi projection with Fermi energy in a spectral gap, its integral kernel is exponentially localized, see (3.1), and $C(P)$ is connected to the quantum transport coefficients of the physical system modelled by $H_{b}$. In particular, the Hall conductivity of the system computed in the linear response regime is proportional to $C(P)$ (see for example the recent papers [21, 31] and references therein). For this reason, these materials are also called Chern insulators. Since the integer quantum Hall effect is mainly a two-dimensional phenomenon (even though other types of topological insulators exist in all dimensions), we restrict our analysis to magnetic Hamiltonians in $L^{2}({\mathbb{R}}^{2})$.
@@
As a by-product, in Section 4 we show that the spectral projections of magnetic Scrödinger operators are always continuous with respect to the magnetic field in the strong operator topology, while in Section 5 we show that they fail to be continuous in the uniform topology at least when they are topologically non-trivial, that is whenever their Chern character is not zero. An explicit example for such a failure is given by the projection on the ground state of the purely magnetic Landau Hamiltonian.
@@
    The resolvent $(H_{b}-z{\bf 1})^{-1}$ is an integral operator with an integral kernel that is jointly continuous outside of the diagonal and such that for every $K_{b}$ compact subset of the resolvent set of $H_{b}$ there exists $C,\alpha>0$ independent of $0\leq b\leq 1$ such that for all ${\bf x}\neq{\bf x}^{\prime}\in{\mathbb{R}}^{2}$:
    $\sup_{z\in K_{b}}\left|(H_{b}-z{\bf 1})^{-1}({\bf x};{\bf x}^{\prime})\right|\leq C\ln\left(2+\|{\bf x}-{\bf x}^{\prime}\|^{-1}\right)e^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}\,.$ (1.5)
@@
    $\sup_{z\in K_{b}}\left|(-\mathrm{i}\nabla_{\bf x}-bA({\bf x}))(H_{b}-z{{\bf 1}})^{-1}({\bf x};{\bf x}^{\prime})\right|\leq C\|{\bf x}-{\bf x}^{\prime}\|^{-1}e^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}\,.$ (1.6)
@@ 2. The quasi-inverse and stability of spectral gaps
In this section we review the basics of gauge covariant magnetic perturbation theory, the presentation is based on [13, 14, 37, 17]. Such a perturbation theory is inspired by the well-known Peierls substitution and it is based on the gauge invariance of the electromagnetic theory. An heuristic discussion of the theory is postponed to Section 2.1.
@@
${A}_{\bf y}({\bf x}):=\left(\int_{0}^{1}s\,\mathfrak{b}({\bf y}+s({\bf x}-{\bf y}))\mathrm{d}s\right)\;(-x_{2}+y_{2},x_{1}-y_{1}).$ (2.1)
@@
$A_{\mathbf{0}}(\vec{x})=\left(\int_{0}^{1}s\mathfrak{b}(s\vec{x})\mathrm{d}s\right)\wedge\vec{x}.$ (2.2)
@@
$\|\partial_{\bf x}^{\alpha}{A}_{\bf y}({\bf x})\|\leq b_{\alpha}\;\|{\bf x}-{\bf y}\|,\quad\alpha\in\mathbb{N}^{2},\quad|\alpha|\leq 1.$ (2.3)
@@
×$\nabla\times{A}_{\bf y}({\bf x})=\nabla\times{A}({\bf x})=\mathfrak{b}({\bf x})\,.$
@@
${A}({\bf x})-{A}_{\bf y}({\bf x})=\nabla_{\bf x}\varphi({\bf x},{\bf y})$ (2.4)
and $\varphi$ is given by the line integral of the vector potential ${A}$ along the straight-line segment connecting ${\bf y}$ and ${\bf x}$
$\varphi({\bf x},{\bf y}):=\int_{{\bf y}}^{{\bf x}}{A}\cdot\mathrm{d}\mathbf{s}\equiv\int_{0}^{1}\,{A}({\bf y}+s({\bf x}-{\bf y}))\cdot({\bf x}-{\bf y})\,\mathrm{d}s.$ (2.5)
@@
$\varphi({\bf x},{\bf y})=-\left(\int_{0}^{1}\int_{0}^{1}\,t\mathfrak{b}(t({\bf y}+s({\bf x}-{\bf y})))\mathrm{d}t\mathrm{d}s\right)(x_{1}y_{2}-x_{2}y_{1})\,.$ (2.6)
@@
$\varphi({\bf x},{\bf y})=\int_{\gamma\left({\bf x},{\bf y}\right)}{A}(\mathbf{s})\cdot\mathrm{d}\mathbf{s}=\int_{{\bf y}}^{{\bf x}}{A}\cdot\mathrm{d}\mathbf{s}\,.$ (2.7)
@@
    Let $\langle{\bf x},{\bf y},\,{\bf x}^{\prime}\rangle$ be the oriented triangle with vertices in ${\bf x},{\bf y}$ and ${\bf x}^{\prime}$ and define the flux of $\mathfrak{b}(x)$ through this triangle by $\mathrm{fl}({\bf x},{\bf y},{\bf x}^{\prime})=\int_{\langle{\bf x},{\bf y},\,{\bf x}^{\prime}\rangle}\;\mathfrak{b}(\tilde{{\bf x}})\cdot\mathrm{d}\sigma(\tilde{{\bf x}})$. Then there exists a constant $C<\infty$ such that
    $\varphi({\bf x},{\bf y})+\varphi({\bf y},{\bf x}^{\prime})-\varphi({\bf x},{\bf x}^{\prime})=\mathrm{fl}({\bf x},{\bf y},{\bf x}^{\prime})\,,$ (2.8)
@@
$\big{(}-\mathrm{i}\nabla_{\bf x}-b{A}({\bf x})\big{)}e^{\mathrm{i}b\varphi({\bf x},{\bf y})}=e^{\mathrm{i}b\varphi({\bf x},{\bf y})}\big{(}-\mathrm{i}\nabla_{\bf x}-b{A}_{\bf y}({\bf x})\big{)}\,.$ (2.9)
@@
$S_{\epsilon}(z)({\bf x};{\bf x}^{\prime}):=e^{\mathrm{i}\epsilon\varphi({\bf x},{\bf x}^{\prime})}(H_{b_{0}}-z{\bf 1})^{-1}({\bf x};{\bf x}^{\prime}).$ (2.10)
@@
$|S_{\epsilon}(z)({\bf x};{\bf x}^{\prime})|$ is pointwise dominated by the integral kernel of the resolvent of $H_{b_{0}}$ which due to (1.5) is dominated by an $L^{1}({\mathbb{R}}^{2})$-function of ${\bf x}-{\bf x}^{\prime}$, thus the Schur test [26, Theorem 5.2] implies that $S_{\epsilon}(z)$ is a bounded operator. ∎
@@
$\big{(}H_{b_{0}+\epsilon}-z{\bf 1}\big{)}\,S_{\epsilon}(z)={{\bf 1}}+T_{\epsilon}(z)$ (2.11)
@@
$\big{(}(-\mathrm{i}\nabla_{\bf x}-b_{0}\,{A}({\bf x}))^{2}+V({\bf x})-\bar{z}{\bf 1}\big{)}\,(H_{b_{0}}-z)^{-1}({\bf x};{\bf y})=\delta({\bf x}-{\bf y})\,.$
@@
$\langle(H_{b}-\bar{z}{\bf 1})\psi,S_{\epsilon}(z)\phi\rangle=\langle\psi,\big{(}\mathbf{1}+T_{\epsilon}(z)\big{)}\phi\rangle.$ (2.12)
From the definition of the integral kernel of $T_{\epsilon}(z)$ we deduce that
$|T_{\epsilon}(z)({\bf x};{\bf x}^{\prime})|\leq C\,|\epsilon|\,e^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}.$ (2.13)
The integral kernel is pointwise dominated by an $L^{1}$ function of ${\bf x}-{\bf x}^{\prime}$, hence by a Schur test [26, Theorem 5.2] we get that $\|T_{\epsilon}(z)\|\leq C\,|\epsilon|$. Since $H_{b}$ is essentially selfadjoint on $C^{\infty}_{0}({\mathbb{R}}^{2})$ and $S_{\epsilon}(z)$ and $T_{\epsilon}(z)$ are bounded operators, we can extend (2.12) to any function $\psi\in\mathrm{D}(H_{b})$ and any function $\phi\in L^{2}({\mathbb{R}}^{2})$. Since $H_{b}$ is selfadjoint, this implies that $S_{\epsilon}(z)$ maps $L^{2}({\mathbb{R}}^{2})$ into the domain of $H_{b}$ and (2.11) is proved.
@@
$\big{(}H_{b}-(\lambda+\mathrm{i}\delta){\bf 1}\big{)}^{-1}=S_{\epsilon}(\lambda+\mathrm{i}\delta)\big{(}\mathbf{1}+T_{\epsilon}(\lambda+\mathrm{i}\delta)\big{)}^{-1}.$ (2.14)
@@
$H_{b}-\lambda{\bf 1}=\Big{(}{\bf 1}+\delta\big{(}H_{b}-(\lambda+\mathrm{i}\delta){\bf 1}\big{)}^{-1}\Big{)}\,\big{(}H_{b}-(\lambda+\mathrm{i}\delta){\bf 1}\big{)}$
@@
$\sup_{z\in K}|\{(H_{b}-z{\bf 1})^{-1}T_{\epsilon}(z)\}({\bf x};{\bf y})|\leq C\,|\epsilon|\,e^{-\alpha\|{\bf x}-{\bf y}\|}\,$
@@
$\sup_{z\in K}\left|(H_{b}-z{\bf 1})^{-1}({\bf x};{\bf y})-S_{\epsilon}(z)({\bf x};{\bf y})\right|\leq C\,|\epsilon|\,e^{-\alpha\|{\bf x}-{\bf y}\|}\,.$ (2.16)
@@
One may show that the edges of spectral islands of magnetic Schrödinger operators are actually Lipshitz in the magnetic field:
@@
$\varphi({\bf x},{\bf y})+\varphi({\bf y},{\bf x}^{\prime})-\varphi({\bf x},{\bf x}^{\prime})=\varphi({\bf x}-{\bf y},{\bf y}-{\bf x}^{\prime})=\mathrm{fl}({\bf x},{\bf y},{\bf x}^{\prime})\,.$ (2.18)

### v2

@@ 1.1. Long-range magnetic perturbations
We begin by describing two simple situations that highlight the singular character of magnetic perturbations. In dimension $d=2$, the free Laplacian $\Delta$ is a selfadjoint operator on the Sobolev space $H^{2}({\mathbb{R}}^{2})$ and its spectrum is purely absolutely continuous and coincides with $0,\infty)$. When we allow a non-zero constant magnetic field oriented perpendicular to the plane, we have to consider the operator $(-\mathrm{i}\nabla-bA_{L})^{2}$, $b>0$, where $A_{L}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ is the magnetic vector potential in the so-called symmetric gauge $A_{L}({\bf x})=\frac{1}{2}(-x_{2},x_{1})$. This is the well-known Landau Hamiltonian, which is selfadjoint on the magnetic Sobolev space $H^{2}_{A_{L}}({\mathbb{R}}^{2})$, given by
(1.1) $H^{2}_{A_{L}}({\mathbb{R}}^{2})=\overline{\left\{f\in C^{\infty}_{0}({\mathbb{R}}^{2})\,|\,(-\mathrm{i}\nabla-bA_{L})^{2}f\in L^{2}({\mathbb{R}}^{2})\right\}}^{\|\cdot\|_{D_{A_{L}}}}$
@@
whose coefficients grow polynomially at infinity. This strongly suggests that the “magnetic perturbation” $W_{b}$ is neither form, nor operator bounded with respect to the free Laplacian. In general it can be proved that the perturbed magnetic Hamiltonian fails to converge in the norm resolvent sense to the unperturbed non-magnetic Hamiltonian when the constant magnetic field goes to zero, see [2, Theorem 6.3].
It turns out that most of the singular behaviour induced by long-range magnetic fields is contained in some highly oscillating unimodular phase factors. In the early 2000s, G. Nenciu and the first Author developed a general method which deals with these singular phases in a systematic way [18, 39]. This framework is nowadays called gauge covariant magnetic perturbation theory. The main purpose of this paper is to outline the main ideas behind this method, together with some physically and mathematically relevant applications.
@@ 1.2. Mathematical framework and results
(1.2) $\sup_{{\bf x}\in{\mathbb{R}}^{2}}|\partial^{\alpha}b({\bf x})|\leq b\,b_{\alpha},\quad\alpha\in\mathbb{N}^{2},\quad|\alpha|\leq 1,\quad 0\leq b_{\alpha}<\infty,\quad 0\leq b\leq 1.$
Since $V$ is infinitesimally bounded with respect to the free Laplacian [41, Theorem XIII.96] and the magnetic potentials are smooth, by using the diamagnetic inequality [2, Theorem 2.3] and the standard theory of Schrödinger operators we get that $H_{b}$ is essentially selfadjoint on $C^{\infty}_{0}({\mathbb{R}}^{2})$ and its domain of selfadjointness is the magnetic Sobolev space $H^{2}_{A}({\mathbb{R}}^{2})$, see (1.1) for the definition of the second magnetic Sobolev space.
@@
(1.3) $C(P)=\lim_{L\to\infty}\frac{2\pi\mathrm{i}}{4L^{2}}\int_{[-L,L]^{2}}\left(P[[P,X_{1}],[P,X_{2}]]\right)({\bf x};{\bf x})\,\mathrm{d}{\bf x}$
exists and it is an integer. The quantity $C(P)$ is called the Chern character of the projection $P$, which can also be interpreted as the trace per unit surface of the double commutator $\mathrm{i}\left(P[[P,X_{1}],[P,X_{2}]]\right)$. The name “Chern” originates from the fact that (1.3) coincides with the Chern number of the Bloch bundle generated by $P$, whenever $H_{b}$ commutes with a unitary representation of the group ${\mathbb{Z}}^{2}$ provided by magnetic translations [40]. If $P$ is the Fermi projection with Fermi energy in a spectral gap, its integral kernel is exponentially localized, see (3.1), and $C(P)$ is connected to the quantum transport coefficients of the physical system modelled by $H_{b}$. In particular, the Hall conductivity of the system computed in the linear response regime is proportional to $C(P)$ (see for example the recent papers [22, 32] and references therein). For this reason, these materials are also called Chern insulators. Since the integer quantum Hall effect is mainly a two-dimensional phenomenon (even though other types of topological insulators exist in all dimensions), we restrict our analysis to magnetic Hamiltonians in $L^{2}({\mathbb{R}}^{2})$.
@@
As a by-product, in Section 4 we show that the spectral projections of magnetic Schrödinger operators are always continuous with respect to the magnetic field in the strong operator topology, while in Section 5 we show that they fail to be continuous in the uniform topology at least when they are topologically non-trivial, that is whenever their Chern character is not zero. An explicit example for such a failure is given by the projection on the ground state of the purely magnetic Landau Hamiltonian.
@@
    The resolvent $(H_{b}-z{\bf 1})^{-1}$ is an integral operator with an integral kernel that is jointly continuous outside of the diagonal and such that for every $K_{b}$ compact subset of the resolvent set of $H_{b}$ there exist $C,\alpha>0$ independent of $0\leq b\leq 1$ such that for all ${\bf x}\neq{\bf x}^{\prime}\in{\mathbb{R}}^{2}$:
    (1.4) $\sup_{z\in K_{b}}\left|(H_{b}-z{\bf 1})^{-1}({\bf x};{\bf x}^{\prime})\right|\leq C\ln\left(2+\|{\bf x}-{\bf x}^{\prime}\|^{-1}\right)e^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}\,.$
@@
    (1.5) $\sup_{z\in K_{b}}\left|(-\mathrm{i}\nabla_{\bf x}-bA({\bf x}))(H_{b}-z{{\bf 1}})^{-1}({\bf x};{\bf x}^{\prime})\right|\leq C\|{\bf x}-{\bf x}^{\prime}\|^{-1}e^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}\,.$
@@ 2. The quasi-inverse and stability of spectral gaps
In this section we review the basics of gauge covariant magnetic perturbation theory, the presentation is based on [18, 19, 39, 15]. Such a perturbation theory is inspired by the well-known Peierls substitution and it relies on the gauge invariance of the electromagnetic theory. A heuristic discussion of the theory is postponed to Section 2.1.
@@
(2.1) ${A}_{\bf y}({\bf x}):=\left(\int_{0}^{1}s\,\mathfrak{b}({\bf y}+s({\bf x}-{\bf y}))\mathrm{d}s\right)\;(-x_{2}+y_{2},x_{1}-y_{1}).$
@@
$A_{\mathbf{0}}(\vec{x})=\left(\int_{0}^{1}s\mathfrak{b}(s\vec{x})\mathrm{d}s\right)\wedge\vec{x}.$
@@
$\|\partial_{\bf x}^{\alpha}{A}_{\bf y}({\bf x})\|\leq b_{\alpha}\;\|{\bf x}-{\bf y}\|,\quad\alpha\in\mathbb{N}^{2},\quad|\alpha|\leq 1.$
@@
$\nabla\times{A}_{\bf y}({\bf x})=\nabla\times{A}({\bf x})=\mathfrak{b}({\bf x})\,.$
@@
(2.2) ${A}({\bf x})-{A}_{\bf y}({\bf x})=\nabla_{\bf x}\varphi({\bf x},{\bf y})$
and $\varphi$ is given by the line integral of the vector potential A{A} along the straight-line segment connecting $\bf y$} and $\bf x$}
(2.3) $\varphi({\bf x},{\bf y}):=\int_{{\bf y}}^{{\bf x}}{A}\cdot\mathrm{d}\mathbf{s}\equiv\int_{0}^{1}\,{A}({\bf y}+s({\bf x}-{\bf y}))\cdot({\bf x}-{\bf y})\,\mathrm{d}s.$
@@
(2.4) $\varphi({\bf x},{\bf y})=-\left(\int_{0}^{1}\int_{0}^{1}\,t\mathfrak{b}(t({\bf y}+s({\bf x}-{\bf y})))\mathrm{d}t\mathrm{d}s\right)(x_{1}y_{2}-x_{2}y_{1})\,.$
@@
(2.5) $\varphi({\bf x},{\bf y})=\int_{\gamma\left({\bf x},{\bf y}\right)}{A}(\mathbf{s})\cdot\mathrm{d}\mathbf{s}=\int_{{\bf y}}^{{\bf x}}{A}\cdot\mathrm{d}\mathbf{s}\,.$
@@
[…]
    (2.6) $\varphi({\bf x},{\bf y})+\varphi({\bf y},{\bf x}^{\prime})-\varphi({\bf x},{\bf x}^{\prime})=\mathrm{fl}({\bf x},{\bf y},{\bf x}^{\prime})\,$,
@@
(2.7) $\big(-\mathrm{i}\nabla_{\bf x}-b{A}({\bf x})\big)e^{\mathrm{i}b\varphi({\bf x},{\bf y})}=e^{\mathrm{i}b\varphi({\bf x},{\bf y})}\big(-\mathrm{i}\nabla_{\bf x}-b{A}_{\bf y}({\bf x})\big)\,.$
@@
(2.8) $S_{\epsilon}(z)({\bf x};{\bf x}^{\prime}):=e^{\mathrm{i}\epsilon\varphi({\bf x},{\bf x}^{\prime})}(H_{b_{0}}-z{\bf 1})^{-1}({\bf x};{\bf x}^{\prime}).$
@@
$|S_{\epsilon}(z)({\bf x};{\bf x}^{\prime})|$ is pointwise dominated by the integral kernel of the resolvent of $H_{b_{0}}$ which due to (1.4) is dominated by an $L^{1}({\mathbb{R}}^{2})$-function of ${\bf x}-{\bf x}^{\prime}$, thus the Schur test [27, Theorem 5.2] implies that $S_{\epsilon}(z)$ is a bounded operator. ∎
@@
(2.9) $\big(H_{b_{0}+\epsilon}-z{\bf 1}\big)\,S_{\epsilon}(z)={{\bf 1}}+T_{\epsilon}(z)$
@@
$\big((-\mathrm{i}\nabla_{\bf x}-b_{0}\,{A}({\bf x}))^{2}+V({\bf x})-z{\bf 1}\big)\,(H_{b_{0}}-z)^{-1}({\bf x};{\bf y})=\delta({\bf x}-{\bf y})\,.$
@@
(2.10) $\langle(H_{b}-\bar{z}{\bf 1})\psi,S_{\epsilon}(z)\phi\rangle=\langle\psi,\big(\mathbf{1}+T_{\epsilon}(z)\big)\phi\rangle.$
From the definition of the integral kernel of $T_{\epsilon}(z)$ we deduce that there exist $C,\alpha^{\prime}>0$ such that
$|T_{\epsilon}(z)({\bf x};{\bf x}^{\prime})|\leq C\,|\epsilon|\,e^{-\alpha^{\prime}\|{\bf x}-{\bf x}^{\prime}\|}.$
The integral kernel is pointwise dominated by an $L^{1}$ function of ${\bf x}-{\bf x}^{\prime}$, hence by a Schur test [27, Theorem 5.2] we get that $\|T_{\epsilon}(z)\|\leq C\,|\epsilon|$. Since $H_{b}$ is essentially selfadjoint on $C^{\infty}_{0}({\mathbb{R}}^{2})$ and $S_{\epsilon}(z)$ and $T_{\epsilon}(z)$ are bounded operators, we can extend (2.10) to any function $\psi\in\mathrm{D}(H_{b})$ and any function $\phi\in L^{2}({\mathbb{R}}^{2})$. Since $H_{b}$ is selfadjoint, this implies that $S_{\epsilon}(z)$ maps $L^{2}({\mathbb{R}}^{2})$ into the domain of $H_{b}$ and (2.9) is proved.
@@
(2.11) $\big(H_{b}-(\lambda+\mathrm{i}\delta){\bf 1}\big)^{-1}=S_{\epsilon}(\lambda+\mathrm{i}\delta)\big(\mathbf{1}+T_{\epsilon}(\lambda+\mathrm{i}\delta)\big)^{-1}.$
@@
$H_{b}-\lambda{\bf 1}=\Big({\bf 1}+\mathrm{i}\delta\big(H_{b}-(\lambda+\mathrm{i}\delta){\bf 1}\big)^{-1}\Big)\,\big(H_{b}-(\lambda+\mathrm{i}\delta){\bf 1}\big)$
@@
$\sup_{z\in K}|\{(H_{b}-z{\bf 1})^{-1}T_{\epsilon}(z)\}({\bf x};{\bf y})|\leq C\,|\epsilon|\,e^{-\alpha^{\prime}\|{\bf x}-{\bf y}\|}\,$
@@
(2.13) $\sup_{z\in K}\left|(H_{b}-z{\bf 1})^{-1}({\bf x};{\bf y})-S_{\epsilon}(z)({\bf x};{\bf y})\right|\leq C\,|\epsilon|\,e^{-\alpha^{\prime}\|{\bf x}-{\bf y}\|}\,.$
@@
One may show that the edges of spectral islands of magnetic Schrödinger operators are actually Lipschitz in the magnetic field:
@@
(2.15) $\varphi({\bf x},{\bf y})+\varphi({\bf y},{\bf x}^{\prime})-\varphi({\bf x},{\bf x}^{\prime})=\varphi({\bf x}-{\bf y},{\bf y}-{\bf x}^{\prime})=\mathrm{fl}({\bf x},{\bf y},{\bf x}^{\prime})\,.$